Statistics from a county public health nurse is provided below. Assist her in completing her report by determining the rate for each vital statistic based on the following information.

if a rate can not be calculated, note what information is missing. Provide the rate or information needed in column J, coinciding with the information on that rate. For this county, the midyear population is 500,000, the estimated female population at that time was 75,000.

*Note that all rates are multiplied by 1,000 except Maternal Mortality is by a factor of 10,000.

Neonatal deaths: (<28 days) 18
Answer #1

Total number of population - 500000

Never of female population is 75000

Number of neonatal death -18

The missing data is number of live births ,if it was available then we can find out the neonatal mortality rate .

We can find out the crude death rate of neonates by the given data .

Crude death rate = number of death of neonates /total population ×1000

= 18/500000×1000

= 18/500= 0.036